OpenStudy (anonymous):

Convert to rectangular coordinates: r+rcos(theta)=3

terenzreignz (terenzreignz):

Use thees relations... \[\huge r=\sqrt{x^2+y^2}\] and \[\huge \cos(\theta) = \frac{x}r\] This particular one isn't relevant here, but it might be useful somewhere else in your coursework... \[\huge \color{blue}{\sin(\theta) = \frac{y}{r}}\]

terenzreignz (terenzreignz):

rats... Use *these* relations

OpenStudy (anonymous):

I know those relations. I just don't know how to solve this problem.

terenzreignz (terenzreignz):

Okay... \[\huge \color{red}r +r\color{green}{\cos(\theta)} = 3\] And replace accordingly... \[\huge \color{red}{r=\sqrt{x^2+y^2}}\] \[\huge \color{green}{\cos(\theta) = \frac{x}r}\] Or more directly... \[\huge r\color{green}{\cos(\theta)}=\color{black}x\]

OpenStudy (anonymous):

okay thanks I got it

OpenStudy (anonymous):

btw your responses are oh so nice and colorful

terenzreignz (terenzreignz):

I love emphasis :)